Ask (a) A from the sunlit of the tent. Calculate the temperature of the tent canvas (in °C) based on the following information: The shirtless riders skin temperature is 34.0°C and has an emissivity of 0.965. The exposed area of skin is 0.320 m2. He receives radiation at the r.te of 18.0 W-half what you would calculate if the entire region behind him was hot. The rest of the surroundings are at 34.0°C. eC if the sunlit side of the tent was nearly pure white and if the rider was covered by a white tunic. lReading 4/15/2018
Answer #1

(a)

Given that,

emmisivity, e = 0.965

exposed area, A = 0.32 m^2

rate of heat flow, Q/t = 18 W

T1 = 34 oC = 307 K

Using stefen - boltzmann law of radiation,

Q / t = \sigmae*(A / 2) * (T2^4 - T1^4)

T2 = [((2*Q / t) / \sigmaeA) + T1^4]1/4

Here, \sigma = 5.67*10-8  J/s . m^2 . K^4

T2 = [(2*18 / 5.67*10-8 * 0.965 * 0.32) + 3074] 1/4

T2 = 323.4 K

T2 = 50.4 oC

(b)

lf sunlite side of tent was nearly pure white and rider was covered by a white tunic,

As we know that, white obects more reflect radiations.

Hence, white side of tent and white tunic of rider would prevent the sunlight to enter inside.

so the temperature of the tent canvas would be less than 50.4 oC